mathslover (mathslover):

if : \[\large{\frac{a}{b}=\frac{c}{d}=\frac{e}{f}}\] prove that: \[\large{\frac{2a^4b^2+3a^2e^2-5e^4f}{2b^6+3b^2f^2-5f^5}=\frac{a^4}{b^4}}\]

mathslover (mathslover):

i started like this: \[\large{\frac{a}{b}+\frac{c}{d}+\frac{e}{f}=k}\]

mathslover (mathslover):

\[\large{a=bk ,\quad c=dk,\quad e=fk}\]

mathslover (mathslover):

i put the values of a and e there but still not getting what i wanted

OpenStudy (anonymous):

exactly with ur notations replace \(a^4\) , \(a^2\) , \(e^2\) and \(e^4\) in numerator

OpenStudy (anonymous):

what will u get?

mathslover (mathslover):

yes i did : wait lemme show u what i got

mathslover (mathslover):

\[\large{\frac{2b^6k^4+3b^2f^2k^4-5f^5k^4}{2b^6+3b^2f^2-5f^5}}\]

mathslover (mathslover):

so the answer is k^4

mathslover (mathslover):

that is : \[\large{\frac{a^4}{b^4}}\]

mathslover (mathslover):

right?

OpenStudy (anonymous):

u got it

mathslover (mathslover):

oh ... i got the answer yay !!!thanks a lot @mukushla i dont know what the "hell" i was doing then? thanks again

mathslover (mathslover):

sorry all for disturbing u ..

OpenStudy (anonymous):

np....:)